An observer stands at the bottom of a freeway off-ramp and records the turning direction (L=left, R=right) of each of three successive vehicles. Consider the following events: A = event that exactly one car turns left B = event that all cars turn in the same direction How many outcomes does the event A union B have?
Answer #1

number of events in A N(A)=3   (LRR , RLR , RRL)

number of events in B N(B)=2 (LLL, RRR)

number of events in both N(A n B)=0

hence outcomes does the event A union B have =N(A u B)=N(A)+N(B)-N(A n B)=3+2-0 =5